OpenStudy (anonymous):

Find the measure of each numbered angle. m<1=5x-2,m<2=2x+8

OpenStudy (anonymous):



OpenStudy (anonymous):

Is it the bottom right and left angles that are given the quantities 5x-2 and 2x+8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The sum of all angles inside a triangle is equal to 180. A+B+C=180 (5x-2)+(2x-8)+C=180 So measure of the last angle would be -7x+170. Do you have any other information? Like, what species the triangle is?

OpenStudy (anonymous):

no it just says find the measure of each measured angle